OpenStudy (idealist10):

Find all solutions of y'-(1+x)/3x*y=y^4.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (idealist10):

v=y^-3=1/y^3 y^3=1/v y=(1/v)^(1/3) (1/3)(1/v)^(-2/3)*dv/dx-(1+x)/3x*(1/v)^(1/3)=(1/v)^(4/3) (v^2/3)dv/dx-(1+x)/3x*(1/v)=(1/v)^4 dv/dx-(1+x)/3x*3/v^3=(1/v^4)(3/v^2) dv/dx-(1+x)/(xv^3)=3/v^6 u(x)=e^(-ln abs(x)-x)=1/x*e^-x=1/xe^x

OpenStudy (anonymous):

Is the second term \(\dfrac{1+x}{3x}y\) or \(\dfrac{1+x}{3}xy\) ?

OpenStudy (idealist10):

the first one you wrote.

OpenStudy (anonymous):

\[\begin{align*} y'-\frac{1+x}{3x}y&=y^4\\ y^{-4}y'-\frac{1+x}{3x}y^{-3}&=1\\ -\frac{1}{3}v'-\frac{1+x}{3x}v&=1&\text{letting }v=y^{-3}\text{, so }v'=-3y^{-4}y'\\ v'+\frac{1+x}{x}v&=-3\end{align*}\] Integrating factor: \[\mu(x)=\exp\left(\int \frac{1+x}{x}~dx\right)=\exp\left(\ln|x|+x\right)=xe^x\] \[\begin{align*} xe^xv'+(xe^x)\frac{1+x}{x}v&=-3xe^x\\ xe^xv'+(e^x+xe^x)v&=-3xe^x\\ \frac{d}{dx}\left[xe^xv\right]&=-3xe^x\\ xe^xv&=-3\int xe^x~dx\\ xe^xy^{-3}&=-3\left(xe^x-\int e^x~dx\right)&\text{(integrating by parts)}\\ xe^xy^{-3}&=-3xe^x+3e^x+C\\ y^{-3}&=-3+\frac{3}{x}+\frac{C}{xe^x} \end{align*}\]

OpenStudy (idealist10):

Wait a minute.

OpenStudy (anonymous):

I have to get going, sorry... I should be back on later. I provided an implicit solution, but solving for an explicit one shouldn't be too hard.

OpenStudy (idealist10):

That's fine. I can solve for the explicit solution. Thanks.